# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v1

@@ 6.4.1 Configurations of points
[…]
@@
Note that we require the weights $a_{i}$ to be positive. Informally, we can think of $a_{i}$ as the mass of the point $p_{i}$
@@
Let $\mathbf{a}\in\mathbb{R}^{\mathcal{H}}$ be a weight vector with components $a_{H}>0$ for all $H\in\mathcal{H}$. Let $m=|\mathcal{H}|$ and label the hyperplanes =$\mathcal{H}=\{H_{1},\ldots,H_{m}\}$. Write $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ with $p_{i}=H_{i}^{\perp}$ and $\mathbf{a}=(a_{1},\ldots,a_{m})$ with $a_{i}=a_{H_{i}}$.
@@ 6.4.3 Stability of pairs
[…]
@@
Definition 6.38 is broader than klt condition, in the sense that the strict inequalities $<$ in (1.1) are relaxed to weak inequalities $\leq$ in (6.25).
@@
(i) If $s>n+1$ then the rescaling factor $\lambda=(n+1)/s$ in Equation (6.11) is $<1$. Therefore, the weights $a^{\prime}_{H}=\lambda\cdot a_{H}$ of the rescaled arrangement $(\mathcal{H},\mathbf{a}^{\prime})$ satisfy $a^{\prime}_{H}<a_{H}$ for all $H$. This implies that the weak inequalities $\leq$ in (6.25) become strict inequalities $<$; showing that $(\mathcal{H},\mathbf{a}^{\prime})$ is klt.
@@ 7.1 Critical points of $Q$ and induced arrangements
We use the notation for deletion and restriction triples as in [OT92, Definition 1.14]. Fix $H\in\mathcal{H}$ and write $(\mathcal{H},\mathcal{H}^{\prime},\mathcal{H}^{\prime\prime})$ where =$\mathcal{H}^{\prime}=\mathcal{H}\setminus\{H\}$ and ′′$\mathcal{H}^{\prime\prime}=\mathcal{H}^{H}$ is the induced arrangement obtained by intersecting $H$ with members of $\mathcal{H}^{\prime}$ as defined in Section 2.1 .
@@
[…]
@@ 7.2 Hirzebruch arrangements
In Lemma 7.18 we require that $\mathcal{H}$ is essential. If $\mathcal{H}\subset\mathbb{CP}^{3}$ is an arrangement of $4=2k$ planes that intersect at a common point and whose pairwise intersection are all distinct, then every member of $\mathcal{H}$ intersects the others along $13=k+$ 1 lines but $\mathcal{H}$ is not Hirzebruch.
@@ 7.4.1 Matroids: basic definitions
In a nutshell, matroids are combinatorial structures that mimic finite collections of vectors in a vector space, including information about their linear dependencies (see Example 7.31). Formally, a matroid is a pair $M=(E,\mathcal{B})$, where $E$ is a finite set and $\mathcal{B}$ is a non-empty collection of subsets $B\subset E$ which satisfy the exchange property: for any $B_{1},B_{2}\in\mathcal{B}$ and $b_{1}\in B_{1}\setminus B_{2}$ there exists $b_{2}\in B_{2}\setminus B_{1}$ such that $\left(B_{1}\setminus\{b_{1}\}\right)\cup\{b_{2}\}$ is in $\mathcal{B}$. The elements of $\mathcal{B}$ are called basis of the matroid $M$. The set $E$ is referred as the ground set of $M$. Two matroids $M_{1}=(E_{1},\mathcal{B}_{1})$ and $M_{2}=(E_{2},\mathcal{B}_{2})$ are isomorphic if there is a bijection $f:E_{1}\to E_{2}$ such that $B_{1}\in\mathcal{B}_{1}$ if and only if $B_{2}=f(B_{1})\in\mathcal{B}_{2}$
@@
Let =$\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be a finite collection of pairwise distinct hyperplanes in $\mathbb{P}_{k}^{n}=\mathbb{P}(k^{n+1})$ with empty common intersection. For each $H_{i}$ choose a defining linear equation $h_{i}\in(k^{n+1})^{*}$. The collection of vectors $E=\{h_{1},\ldots,h_{N}\}$ defines, as in Example 7.31 , a matroid $M=(E,\mathcal{B})$. The elements of $\mathcal{B}$ correspond to $(n+1)$-tuples of hyperplanes with empty common intersection. The fact that each $h_{i}\neq 0$ and that every pair of hyperplanes $H_{i}$ and $H_{j}$ with $i\neq j$ satisfy $H_{i}\neq H_{j}$, implies that $M$ is simple.
The correspondence $L\mapsto\mathcal{H}_{L}$ (where $L$ is a linear subspace and $\mathcal{H}_{L}$ is the set of hyperplanes in $\mathcal{H}$ that contain it) defines an isomorphism between the intersection poset $\overline{\mathcal{L}}$ of $\mathcal{H}$ equipped with the order by reverse inclusion and the poset of flats $\mathcal{F}$ equipped with the order by inclusion. The whole space $\mathbb{P}_{k}^{n}\in\overline{\mathcal{L}}$ (corresponding to the intersection of an empty collection of hyperplanes) is mapped to the minimal element $\hat{0}=\overline{\emptyset}$ of $\mathcal{F}$, while $\emptyset\in\overline{\mathcal{L}}$ (the centre of the arrangement) corresponds to the unique maximal flat $\hat{1}$ (the join of all $h_{i}$).
@@ 7.4.2 Quadratic form for matroids
Let $M$ be a simple matroid of rank $n+1$ on the ground set $[N]=\{1,\ldots,N\}$. Let $F$ be a rank two flat. We say that $F$ is irreducible if $|F|\geq 3$. Otherwise, if $|F|=2$ then $F$ is reducible.
@@ 7.4.3 Examples
The Non-Pappus matroid $M$ is the matroid of rank $3$ on the set of nine elements $[9]=\{1,\ldots,9\}$. This matroid has $8$ circuits (minimal dependent sets) shown in Figure 3 as lines connecting triplets of dependant points. The number of bases is equal to
@@
The matroid $M$ is not representable over any field. However, the statement of Theorem 7.39 holds for $M$. Indeed, it can be checked with computer [^7] See https://github.com/MdeBorbon/nonpappus that the matrix $P=-Q$ is copositive, meaning that $\mathbf{x}^{t}\cdot P\cdot\mathbf{x}\geq 0$ for any vector $\mathbf{x}\in\mathbb{R}^{9}$ with components $x_{i}\geq 0$.
@@ A.1.1 Filtrations of vector spaces
A filtration =$\mathcal{F}=\{F_{a}\,|\,a\in\mathbb{R}\}$ of $V$ is a family of vector subspaces $F_{a}\subset V$ parametrized by $a\in\mathbb{R}$ satisfying the following conditions.
@@
Given a filtration =$\mathcal{F}=\{F_{a}\}$ of $V$, we write
@@ A.1.2 Tuples of filtrations and splittings
The set $\operatorname{wt}(\mathcal{F})$ is contained in $\operatorname{wt}(\mathcal{F}^{1})\times\ldots\times\operatorname{wt}(\mathcal{F}^{k})$. In particular, $\operatorname{wt}(\mathcal{F})$ is a finite subset of $[0,1]^{k}$.
@@
    There is a basis $B$ of $U$ such that $B_{i}=B\cap U_{i}$ is a basis of $U_{i}$ , the sets $B_{i}$ are pairwise disjoint, and $B=\cup_{i}B_{i}$
@@
If $k=1$ and =$\mathcal{F}=\{F_{a}\,|\,a\in\mathbb{R}\}$, then we can construct a splitting of $\mathcal{F}$ as follows. Set $U_{a}=\{0\}$ if $a\notin\operatorname{wt}(\mathcal{F})$ and, for every $a\in\operatorname{wt}(\mathcal{F})$ choose a linear subspace $U_{a}\subset F_{a}$ such that $F_{a}=F_{<a}\oplus U_{a}$. It is then easy to verify that $\{U_{a}\,|\,a\in\mathbb{R}\}$ satisfies (A.3) and (A.4) . Note that the subspaces $U_{a}$ are isomorphic to the graded components of $\operatorname{Gr}_{a}=F_{a}\,/\,F_{<a}$ of $\mathcal{F}$.
@@
the maps $\Phi_{\mathbf{a}}$ define a linear isomorphism $\Phi\in GL(V)$ with $\Phi(U_{\mathbf{a}})=\widetilde{U}_{\mathbf{a}}$ If follows from Equation (A.4) , that $\Phi$ preserves each of the subspaces $F^{i}_{a}$ of the filtration $\mathcal{F}^{i}$ for all $i$, so $\Phi(\mathcal{F})=\mathcal{F}$. ∎
@@
If $\mathbf{a}\in\operatorname{wt}(\mathcal{F})$ then $F_{\mathbf{a}}=L_{i}$, where $i$ is the unique component of $\mathbf{a}$ with $a_{i}=0$, and $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}=0$ Therefore, if Equation (A.6) is satisfied, we must take $U_{\mathbf{a}}=L_{i}$. The subspaces $U_{\mathbf{a}}$ do not make a direct sum. Therefore, the filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are not compatible.
@@ A.1.3 Adapted bases
(i)$\implies$(ii). Suppose that $\{U_{\mathbf{a}}\,|\,\mathbf{a}\in\mathbb{R}^{k}\}$ is a splitting of $\mathcal{F}$. For each $\mathbf{a}\in\mathbb{R}^{k}$ such that $U_{\mathbf{a}}$ is non-zero, choose a basis $B_{\mathbf{a}}$ of $U_{\mathbf{a}}$. Then $B=\bigcup_{\mathbf{a}}B_{\mathbf{a}}$ is adapted to $\mathcal{F}$.
(ii)$\implies$(i). Suppose that $B$ is a basis of $V$ adapted to $\mathcal{F}$. We want to define subspaces $U_{\mathbf{a}}\subset V$ that split $\mathcal{F}$. To do this, for each $e\in B$ we consider the subspace $F_{\mathbf{a}}$ with smallest $\mathbf{a}$ that contains $e$. In detail, note that if both $F_{\mathbf{a}}$ and $F_{\mathbf{a}^{\prime}}$ contain $e$, then $F_{\min\{\mathbf{a},\mathbf{a}^{\prime}\}}$ also does, where $\min\{\mathbf{a},\mathbf{a}^{\prime}\}$ is the vector with components $\min\{a_{i},a^{\prime}_{i}\}$. Thus, we have a map $\Phi:B\to\mathbb{R}^{k}$ given by
@@
If $k=2$ then any two filtrations $\mathcal{F}^{1}$ and $\mathcal{F}^{2}$ on $V$ are compatible, see [Xu24, Lemma 3.5]. However, if $k\geq 3$ then generic $k$-tuples of filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are not compatible. For example if $k=3$, $n=2$ and $\mathcal{F}^{i}$ define $3$ distinct lines in $\mathbb{C}^{2}$ then it is clear that there is no basis of $\mathbb{C}^{2}$ adapted to $\mathcal{F}^{1},\mathcal{F}^{2},\mathcal{F}^{3}$; c.f. Example A.13.
@@ A.1.4 Nested sets of linear subspaces
We proceed by induction on the dimension $n$ of the ambient space $V^{*}$. The statement is trivially true if $n=1$. Let $\mathcal{N}$ be a *-nested set in $V^{*}$ with $n=\dim V^{*}$ and assume that the statement holds true for dimensions $<n$. We want to show that there is a basis $B^{*}$ of $V^{*}$ adapted to $\mathcal{N}$.
@@
On the other hand, using item (ii) in Definition A.26 , it is not hard to show that if $\mathcal{S}$ is a nested set in $V$, then $|\mathcal{S}|\leq\dim V$. If item (ii) is not satisfied then this is no longer true. For example, if $L_{1}$ and $L_{2}$ are two distinct lines in $V=\mathbb{C}^{2}$ going through the origin, then the set $\mathcal{S}=\{\{0\},L_{1},L_{2}\}$ satisfies item (i) of Definition A.26 but it does not satisfy item (ii); and $23=|\mathcal{S}|>\dim V=$ 2.
@@ A.2.1 Multivectors and subspaces
More generally, for any non-zero $v\in\Lambda^{r}V$, the dimension of $\ker(\wedge v)$ is $\leq r$ and it is $=r$ precisely when $v$ is decomposable. The set of decomposable vectors in $\Lambda^{r}V$ is the zero locus of a set of homogeneous quadratic equations, known as the Plücker relations.

### v2

@@ 6.4.1 Configurations of points
of points $p_{i}\in\mathbb{CP}^{n}$ together with a weight vector $\mathbf{a}=(a_{1},\ldots,a_{m})\in\mathbb{R}^{m}_{>0}$.
@@
Note that we require the weights $a_{i}$ to be positive. Informally, we can think of $a_{i}$ as the mass of the point $p_{i}$ .
@@
Let $\mathbf{a}\in\mathbb{R}^{\mathcal{H}}$ be a weight vector with components $a_{H}>0$ for all $H\in\mathcal{H}$. Let $m=|\mathcal{H}|$ and label the hyperplanes $\mathcal{H}=\{H_{1},\ldots,H_{m}\}$. Write $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ with $p_{i}=H_{i}^{\perp}$ and $\mathbf{a}=(a_{1},\ldots,a_{m})$ with $a_{i}=a_{H_{i}}$.
@@ 6.4.3 Stability of pairs
$\begin{gathered}\forall\,H\in\mathcal{H}:\,\,0<a_{H}\leq 1\,,\\ \forall\,L\in\mathcal{L}:\,\,\sum_{H|L\subset H}a_{H}\leq\operatorname{codim}L\,.\end{gathered}$ (6.25)
@@
Definition 6.38 is broader than klt condition, in the sense that the strict inequalities << in (1.1) are relaxed to weak inequalities $\leq$ in (6.25).
@@
(i) If $s>n+1$ then the rescaling factor $\lambda=(n+1)/s$ in Equation (6.11) is <1<1. Therefore, the weights $a^{\prime}_{H}=\lambda\cdot a_{H}$ of the rescaled arrangement $(\mathcal{H},\mathbf{a}^{\prime})$ satisfy $a^{\prime}_{H}<a_{H}$ for all $H$. This implies that the weak inequalities $\leq$ in (6.25) become strict inequalities <<; showing that $(\mathcal{H},\mathbf{a}^{\prime})$ is klt.
@@ 7.1 Critical points of $Q$ and induced arrangements
We use the notation for deletion and restriction triples as in [OT92, Definition 1.14]. Fix $H\in\mathcal{H}$ and write (,,′′)($\mathcal{H}$$\mathcal{H}^{\prime},\mathcal{H}^{\prime\prime}$) where $\mathcal{H}^{\prime}=\mathcal{H}\setminus\{H\}$ and $\mathcal{H}^{\prime\prime}=\mathcal{H}^{H}$ is the induced arrangement obtained by intersecting $H$ with members of $\mathcal{H}^{\prime}$ as defined in Section 2.1 .
@@
$=2n\cdot s^{2}\,-2(n+1)\cdot\sum_{H\in\mathcal{H}}\,\,\sum_{H^{\prime}|\,H^{\prime}\pitchfork H}a_{H}\cdot a_{H^{\prime}}$
@@ 7.2 Hirzebruch arrangements
In Lemma 7.18 we require that $\mathcal{H}$ is essential. If $\mathcal{H}\subset\mathbb{CP}^{3}$ is an arrangement of $4=2k$ planes that intersect at a common point and whose pairwise intersection are all distinct, then every member of $\mathcal{H}$ intersects the others along $3=k+1$ lines but $\mathcal{H}$ is not Hirzebruch.
@@ 7.4.1 Matroids: basic definitions
In a nutshell, matroids are combinatorial structures that mimic finite collections of vectors in a vector space, including information about their linear dependencies (see Example 7.31). Formally, a matroid is a pair $M=(E,\mathcal{B})$, where $E$ is a finite set and $\mathcal{B}$ is a non-empty collection of subsets $B\subset E$ which satisfy the exchange property: for any $B_{1},B_{2}\in\mathcal{B}$ and $b_{1}\in B_{1}\setminus B_{2}$ there exists $b_{2}\in B_{2}\setminus B_{1}$ such that $\left(B_{1}\setminus\{b_{1}\}\right)\cup\{b_{2}\}$ is in $\mathcal{B}$. The elements of $\mathcal{B}$ are called basis of the matroid $M$. The set $E$ is referred as the ground set of $M$. Two matroids $M_{1}=(E_{1},\mathcal{B}_{1})$ and $M_{2}=(E_{2},\mathcal{B}_{2})$ are isomorphic if there is a bijection $f:E_{1}\to E_{2}$ such that $B_{1}\in\mathcal{B}_{1}$ if and only if $B_{2}=f(B_{1})\in\mathcal{B}_{2}$ .
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be a finite collection of pairwise distinct hyperplanes in $\mathbb{P}_{k}^{n}=\mathbb{P}(k^{n+1})$ with empty common intersection. For each $H_{i}$ choose a defining linear equation $h_{i}\in(k^{n+1})^{*}$. The collection of vectors $E=\{h_{1},\ldots,h_{N}\}$ defines, as in Example 7.31 , a matroid $M=(E,\mathcal{B})$. The elements of $\mathcal{B}$ correspond to $(n+1)$-tuples of hyperplanes with empty common intersection. The fact that each $h_{i}\neq 0$ and that every pair of hyperplanes $H_{i}$ and $H_{j}$ with $i\neq j$ satisfy $H_{i}\neq H_{j}$, implies that $M$ is simple.
The correspondence $L\mapsto\mathcal{H}_{L}$ (where $L$ is a linear subspace and $\mathcal{H}_{L}$ is the set of hyperplanes in $\mathcal{H}$ that contain it) defines an isomorphism between the intersection poset $\overline{\mathcal{L}}$ of $\mathcal{H}$ equipped with the order by reverse inclusion and the poset of flats $\mathcal{F}$ equipped with the order by inclusion. The whole space $\mathbb{P}_{k}^{n}\in\overline{\mathcal{L}}$ (corresponding to the intersection of an empty collection of hyperplanes) is mapped to the minimal element $\hat{0}$$\overline{\emptyset}$ of $\mathcal{F}$, while $\emptyset\in\overline{\mathcal{L}}$ (the centre of the arrangement) corresponds to the unique maximal flat $\hat{1}$ (the join of all $h_{i}$).
@@ 7.4.2 Quadratic form for matroids
Let $M$ be a simple matroid of rank $n+1$ on the ground set $N]=\{1,\ldots,N\}$. Let $F$ be a rank two flat. We say that $F$ is irreducible if $|F|\geq 3$. Otherwise, if $|F|=2$ then $F$ is reducible.
@@ 7.4.3 Examples
The Non-Pappus matroid $M$ is the matroid of rank $3$ on the set of nine elements $9]=\{1,\ldots,9\}$. This matroid has $8$ circuits (minimal dependent sets) shown in Figure 3 as lines connecting triplets of dependant points. The number of bases is equal to
@@
The matroid $M$ is not representable over any field. However, the statement of Theorem 7.39 holds for $M$. Indeed, it can be checked with computer [^7] See https://github.com/MdeBorbon/nonpappus that the matrix P=QP=-Q is copositive, meaning that $\mathbf{x}^{t}\cdot P\cdot\mathbf{x}\geq 0$ for any vector $\mathbf{x}\in\mathbb{R}^{9}$ with components $x_{i}\geq 0$.
@@ A.1.1 Filtrations of vector spaces
A filtration $\mathcal{F}=\{F_{a}\,|\,a\in\mathbb{R}\}$ of $V$ is a family of vector subspaces $F_{a}\subset V$ parametrized by $a\in\mathbb{R}$ satisfying the following conditions.
@@
Given a filtration $\mathcal{F}=\{F_{a}\}$ of $V$, we write
@@ A.1.2 Tuples of filtrations and splittings
The set $\operatorname{wt}(\mathcal{F})$ is contained in $\operatorname{wt}(\mathcal{F}^{1})\times\ldots\times\operatorname{wt}(\mathcal{F}^{k})$. In particular, $\operatorname{wt}(\mathcal{F})$ is a finite subset of $0,1]^{k}$.
@@
    There is a basis $B$ of $U$ such that $B_{i}=B\cap U_{i}$ is a basis of $U_{i}$ , the sets $B_{i}$ are pairwise disjoint, and $B=\cup_{i}B_{i}$ .
@@
If $k=1$ and $\mathcal{F}=\{F_{a}\,|\,a\in\mathbb{R}\}$, then we can construct a splitting of $\mathcal{F}$ as follows. Set $U_{a}=\{0\}$ if $a\notin\operatorname{wt}(\mathcal{F})$ and, for every $a\in\operatorname{wt}(\mathcal{F})$ choose a linear subspace $U_{a}\subset F_{a}$ such that $F_{a}=F_{<a}\oplus U_{a}$. It is then easy to verify that $\{U_{a}\,|\,a\in\mathbb{R}\}$ satisfies (A.3) and (A.4) . Note that the subspaces $U_{a}$ are isomorphic to the graded components of $\operatorname{Gr}_{a}=F_{a}\,/\,F_{<a}$ of $\mathcal{F}$.
@@
the maps $\Phi_{\mathbf{a}}$ define a linear isomorphism $\Phi\in GL(V)$ with $\Phi(U_{\mathbf{a}})=\widetilde{U}_{\mathbf{a}}$ . If follows from Equation (A.4) , that $\Phi$ preserves each of the subspaces $F^{i}_{a}$ of the filtration $\mathcal{F}^{i}$ for all $i$, so $\Phi(\mathcal{F})=\mathcal{F}$. ∎
@@
If $\mathbf{a}\in\operatorname{wt}(\mathcal{F})$ then $F_{\mathbf{a}}=L_{i}$, where $i$ is the unique component of $\mathbf{a}$ with $a_{i}=0$, and $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}=0$ . Therefore, if Equation (A.6) is satisfied, we must take $U_{\mathbf{a}}=L_{i}$. The subspaces $U_{\mathbf{a}}$ do not make a direct sum. Therefore, the filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are not compatible.
@@ A.1.3 Adapted bases
[…]
@@
If $k=2$ then any two filtrations $\mathcal{F}^{1}$ and $\mathcal{F}^{2}$ on $V$ are compatible, see [Xu24, Lemma 3.5]. However, if $k\geq 3$ then generic $k$-tuples of filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are not compatible. For example if $k=3$, $n=2$ and $\mathcal{F}^{i}$ define 33 distinct lines in $\mathbb{C}^{2}$ then it is clear that there is no basis of $\mathbb{C}^{2}$ adapted to $\mathcal{F}^{1},\mathcal{F}^{2},\mathcal{F}^{3}$; c.f. Example A.13.
@@ A.1.4 Nested sets of linear subspaces
We proceed by induction on the dimension $n$ of the ambient space $V^{*}$. The statement is trivially true if $n=1$. Let $\mathcal{N}$ be a *-nested set in $V^{*}$ with $n=\dim V^{*}$ and assume that the statement holds true for dimensions <n<n. We want to show that there is a basis $B^{*}$ of $V^{*}$ adapted to $\mathcal{N}$.
@@
On the other hand, using item (ii) in Definition A.26 , it is not hard to show that if $\mathcal{S}$ is a nested set in $V$, then $|\mathcal{S}|\leq\dim V$. If item (ii) is not satisfied then this is no longer true. For example, if $L_{1}$ and $L_{2}$ are two distinct lines in $V=\mathbb{C}^{2}$ going through the origin, then the set $\mathcal{S}=\{\{0\},L_{1},L_{2}\}$ satisfies item (i) of Definition A.26 but it does not satisfy item (ii); and $3=|\mathcal{S}|>\dim V=2$.
@@ A.2.1 Multivectors and subspaces
More generally, for any non-zero $v\in\Lambda^{r}V$, the dimension of $\ker(\wedge v)$ is $\leq r$ and it is =r=r precisely when $v$ is decomposable. The set of decomposable vectors in $\Lambda^{r}V$ is the zero locus of a set of homogeneous quadratic equations, known as the Plücker relations.